What is the next term in the series 1,1,2,3,5,8,13...?
A)18
B)21
C)13
D)9
E)20

Answer by vleith(2825)

every term is the sum of the two previous terms
1+1 = 2
1+2 = 3
2+3 = 5
etc
so the next term is
8+13 = 21